# Spatial Blind Spot: Auditory Motion Perception Deficits in Audio LLMs

### v1

@@ 4.1 Experimental Setup
We evaluate four representative large audio–language models, including Qwen2-Audio-7B [chu2024qwen2], Voxtral-Mini-3B-2507 [liu2025voxtral], Aero-1-Audio [mandel2023aero], and Kimi-Audio-7B [ding2025kimi], on the proposed AMPBench. All models are tested in a zero-shot setting without any fine-tuning or adaptation. Each system receives binaural audio together with the motion-related question following the QA formats introduced in Section 3. Responses are normalized into canonical labels (A–D or TRUE/FALSE) using a rule-based parser, and any unrecognized output is counted as incorrect. Accuracy is averaged over three random seeds and across all task variants, following standard benchmark evaluation.
@@
Task1 Qwen2-Audio-7B [chu2024qwen2] $40.48{\scriptstyle(\pm 0.12)}$ $42.26{\scriptstyle(\pm 0.07)}$ $\mathbf{40.48}{\scriptstyle(\pm 0.11)}$ $\mathbf{42.86}{\scriptstyle(\pm 0.06)}$ $\mathbf{41.52}{\scriptstyle(\pm 0.09)}$
Voxtral-Mini-3B-2507 [liu2025voxtral] $\mathbf{44.64}{\scriptstyle(\pm 0.10)}$ $37.50{\scriptstyle(\pm 0.13)}$ $39.88{\scriptstyle(\pm 0.14)}$ $39.29{\scriptstyle(\pm 0.08)}$ $40.33{\scriptstyle(\pm 0.12)}$
Aero-1-Audio [mandel2023aero] $42.86{\scriptstyle(\pm 0.05)}$ $39.88{\scriptstyle(\pm 0.11)}$ $39.29{\scriptstyle(\pm 0.10)}$ $41.67{\scriptstyle(\pm 0.06)}$ $40.93{\scriptstyle(\pm 0.07)}$
Kimi-Audio-7B [ding2025kimi] $43.45{\scriptstyle(\pm 0.15)}$ $\mathbf{43.45}{\scriptstyle(\pm 0.09)}$ $39.88{\scriptstyle(\pm 0.12)}$ $39.28{\scriptstyle(\pm 0.08)}$ $\mathbf{41.52}{\scriptstyle(\pm 0.11)}$
Task2 Qwen2-Audio-7B [chu2024qwen2] $41.07{\scriptstyle(\pm 0.10)}$ $\mathbf{41.67}{\scriptstyle(\pm 0.06)}$ $\mathbf{44.64}{\scriptstyle(\pm 0.12)}$ $\mathbf{44.64}{\scriptstyle(\pm 0.07)}$ $\mathbf{43.01}{\scriptstyle(\pm 0.09)}$
Voxtral-Mini-3B-2507 [liu2025voxtral] $\mathbf{44.64}{\scriptstyle(\pm 0.14)}$ $39.88{\scriptstyle(\pm 0.10)}$ $41.67{\scriptstyle(\pm 0.08)}$ $38.69{\scriptstyle(\pm 0.13)}$ $41.22{\scriptstyle(\pm 0.12)}$
Aero-1-Audio [mandel2023aero] $42.26{\scriptstyle(\pm 0.11)}$ $39.88{\scriptstyle(\pm 0.14)}$ $39.88{\scriptstyle(\pm 0.13)}$ $41.67{\scriptstyle(\pm 0.07)}$ $40.92{\scriptstyle(\pm 0.09)}$
Kimi-Audio-7B [ding2025kimi] $42.86{\scriptstyle(\pm 0.08)}$ $39.88{\scriptstyle(\pm 0.07)}$ $38.69{\scriptstyle(\pm 0.15)}$ $39.26{\scriptstyle(\pm 0.10)}$ $40.17{\scriptstyle(\pm 0.11)}$
@@
Voxtral-Mini-3B-2507 [liu2025voxtral] $\mathbf{44.05}{\scriptstyle(\pm 0.08)}$ $40.48{\scriptstyle(\pm 0.14)}$ $40.48{\scriptstyle(\pm 0.09)}$ $39.88{\scriptstyle(\pm 0.07)}$ $41.22{\scriptstyle(\pm 0.12)}$
Aero-1-Audio [mandel2023aero] $42.86{\scriptstyle(\pm 0.06)}$ $39.88{\scriptstyle(\pm 0.07)}$ $39.88{\scriptstyle(\pm 0.11)}$ $41.67{\scriptstyle(\pm 0.08)}$ $41.07{\scriptstyle(\pm 0.10)}$
@@ 4.2 Benchmark Results and Analysis
Although we evaluated multiple models under varying signal-to-noise ratio (SNR) conditions, none of them exhibited noticeable performance changes as the SNR increased, accuracy varies by only a few percentage points when moving from clean ($30$ dB) to severely degraded (0 dB) audio. This suggests that current LALMs are largely insensitive to the overall acoustic quality of the waveform. In other words, the accuracy remained nearly constant across clean and noisy environments, suggesting that current large audio-language models are largely insensitive to acoustic quality and fail to utilize spatial information embedded in the waveform. From a perceptual perspective, this stands in stark contrast to human listeners, whose performance degrades noticeably in low-SNR conditions.
Across tasks, we observe that Task 2, which involves variable-speed motion, is slightly easier for most models than Task 1 and Task 3, indicating that some forms of variation may be easier to capture than front–back ambiguities. Model-wise, Qwen2-Audio-7B tends to dominate in Task 2 and Task 3 at higher SNRs, while Voxtral-Mini-3B-2507 occasionally achieves the best scores at 0 dB, reflecting different robustness trade-offs between architectures. Nevertheless, the overall performance remains far from satisfactory in all regimes.
@@
Table 3: MCQ vs T/F accuracy on AMPBench (clean). Each value reports the mean accuracy with its standard deviation across repeated runs. Bold numbers indicate the best-performing model within each task and question format.
@@ 4.3 Prompt Ablation
To examine whether audio–language models genuinely verify auditory motion rather than following linguistic heuristics, we focus on the True/False (T/F) format of AMPBench. Each audio sample is paired with two statements describing a motion trajectory. The True statement corresponds to the ground-truth direction, whereas the False statement is generated by mirroring the trajectory (e.g., left-to-right versus right-to-left) while matching lexical length, tense, and sentiment to eliminate textual shortcuts, as illustrated in Figure 3.
Figure 3: Examples of True and False statements used in the T/F verification task.
@@
Table 4 summarizes the results under the clean condition of AMPBench.
Across all evaluated models, a consistent behavioral pattern appears within the scope of our motion–verification setting. The hit rate on True statements (TPR) is notably low, with the strongest model, Aero-1-Audio, accepting fewer than $21\%$ of True statements. In contrast, the rejection rate for False statements (TNR) is considerably higher, with Kimi-Audio-7B exceeding $84\%$. This imbalance suggests that under our evaluation protocol, models tend to be conservative when deciding whether a statement matches the underlying motion.
@@
Table 4: True/False verification on AMPBench (clean). We report true positive rate (TPR), true negative rate (TNR), and YesBias (%), where lower YesBias indicates reduced tendency toward “Yes” responses. Bold values mark the best performance for each metric.
@@ 5 Conclusion
In our work, we introduce AMPBench, a physically grounded benchmark designed to systematically evaluate auditory motion perception in large audio–language models (LALMs). Through controlled experiments and cross-model comparisons, we reveal a consistent deficiency in spatial reasoning: while current models can often recognize what is heard, they still struggle to perceive where the sound source is and how it moves. This gap appears across architectures of different scales and training regimes, suggesting a structural limitation rather than model-specific behavior.
@@
Beyond serving as a diagnostic benchmark, AMPBench provides a principled foundation for studying spatially grounded auditory intelligence. The ability to track moving sound sources is essential for real-world, embodied agents that must locate, follow, or respond to dynamic auditory events in their environment. AMPBench thus highlights the need for richer training data, architectural elements that explicitly model spatial cues, and learning objectives that reward accurate reasoning about physical motion. We hope this work encourages further exploration toward audio–language systems capable of perceiving the world not only semantically but also spatially, enabling a more embodied form of auditory understanding.
@@ 6 Limitations
While AMPBench provides a controlled and fully reproducible environment for probing auditory motion reasoning, several limitations still constrain its generality. First, the dataset is built upon analytically rendered spatial audio rather than real-world recordings, which means it cannot fully capture complex acoustic phenomena such as room reverberation, partial occlusion by obstacles, or interference from competing sound sources. These factors could meaningfully influence how models perceive motion in practical scenarios. Second, the present design restricts motion to a set of discrete start and end trajectories. More naturalistic motion, including continuously varying paths, curved trajectories, or motion with acceleration, is not yet represented and may introduce qualitatively different reasoning challenges. Third, the benchmark primarily employs synthetic tones and narrow-band signals to isolate spatial cues. While this choice benefits controllability, it leaves open how natural sound categories, environmental events, or speech-like stimuli interact with motion perception mechanisms in LALMs.
Future benchmarks should therefore incorporate real spatial recordings, continuous and diverse motion trajectories, and a broader set of sound categories to provide a more comprehensive evaluation of spatial reasoning capabilities. We hope AMPBench can serve as a foundation and stepping stone toward the next generation of benchmarks for spatially embodied auditory intelligence.

### v2

@@ 4.1 Experimental Setup
Models and Evaluation Protocol.
We evaluate four representative large audio–language models, including Qwen2-Audio-7B chu2024qwen2, Voxtral-Mini-3B-2507 liu2025voxtral, Aero-1-Audio mandel2023aero, and Kimi-Audio-7B ding2025kimi, on the proposed AudioMotionBench. All models are tested in a zero-shot setting without any fine-tuning or adaptation. Each system receives binaural audio together with the motion-related question following the QA formats introduced in Section 3. Responses are normalized into canonical labels (A–D or TRUE/FALSE) using a rule-based parser, and any unrecognized output is counted as incorrect. Accuracy is averaged over three random seeds and across all task variants, following standard benchmark evaluation.
Decoding and Reporting Details.
@@
SCT Qwen2-Audio-7B chu2024qwen2 $40.48{\scriptstyle(\pm 0.12)}$ $42.26{\scriptstyle(\pm 0.07)}$ $\mathbf{40.48}{\scriptstyle(\pm 0.11)}$ $\mathbf{42.86}{\scriptstyle(\pm 0.06)}$ $\mathbf{41.52}{\scriptstyle(\pm 0.09)}$
Voxtral-Mini-3B-2507 liu2025voxtral $\mathbf{44.64}{\scriptstyle(\pm 0.10)}$ $37.50{\scriptstyle(\pm 0.13)}$ $39.88{\scriptstyle(\pm 0.14)}$ $39.29{\scriptstyle(\pm 0.08)}$ $40.33{\scriptstyle(\pm 0.12)}$
Aero-1-Audio mandel2023aero $42.86{\scriptstyle(\pm 0.05)}$ $39.88{\scriptstyle(\pm 0.11)}$ $39.29{\scriptstyle(\pm 0.10)}$ $41.67{\scriptstyle(\pm 0.06)}$ $40.93{\scriptstyle(\pm 0.07)}$
Kimi-Audio-7B ding2025kimi $43.45{\scriptstyle(\pm 0.15)}$ $\mathbf{43.45}{\scriptstyle(\pm 0.09)}$ $39.88{\scriptstyle(\pm 0.12)}$ $39.28{\scriptstyle(\pm 0.08)}$ $\mathbf{41.52}{\scriptstyle(\pm 0.11)}$
SDT Qwen2-Audio-7B chu2024qwen2 $41.07{\scriptstyle(\pm 0.10)}$ $\mathbf{41.67}{\scriptstyle(\pm 0.06)}$ $\mathbf{44.64}{\scriptstyle(\pm 0.12)}$ $\mathbf{44.64}{\scriptstyle(\pm 0.07)}$ $\mathbf{43.01}{\scriptstyle(\pm 0.09)}$
Voxtral-Mini-3B-2507 liu2025voxtral $\mathbf{44.64}{\scriptstyle(\pm 0.14)}$ $39.88{\scriptstyle(\pm 0.10)}$ $41.67{\scriptstyle(\pm 0.08)}$ $38.69{\scriptstyle(\pm 0.13)}$ $41.22{\scriptstyle(\pm 0.12)}$
Aero-1-Audio mandel2023aero $42.26{\scriptstyle(\pm 0.11)}$ $39.88{\scriptstyle(\pm 0.14)}$ $39.88{\scriptstyle(\pm 0.13)}$ $41.67{\scriptstyle(\pm 0.07)}$ $40.92{\scriptstyle(\pm 0.09)}$
Kimi-Audio-7B ding2025kimi $42.86{\scriptstyle(\pm 0.08)}$ $39.88{\scriptstyle(\pm 0.07)}$ $38.69{\scriptstyle(\pm 0.15)}$ $39.26{\scriptstyle(\pm 0.10)}$ $40.17{\scriptstyle(\pm 0.11)}$
@@
Voxtral-Mini-3B-2507 liu2025voxtral $\mathbf{44.05}{\scriptstyle(\pm 0.08)}$ $40.48{\scriptstyle(\pm 0.14)}$ $40.48{\scriptstyle(\pm 0.09)}$ $39.88{\scriptstyle(\pm 0.07)}$ $41.22{\scriptstyle(\pm 0.12)}$
Aero-1-Audio mandel2023aero $42.86{\scriptstyle(\pm 0.06)}$ $39.88{\scriptstyle(\pm 0.07)}$ $39.88{\scriptstyle(\pm 0.11)}$ $41.67{\scriptstyle(\pm 0.08)}$ $41.07{\scriptstyle(\pm 0.10)}$
@@ 4.2 Benchmark Results and Analysis
Effect of Signal-to-Noise Ratio.
Although we evaluated multiple models under varying signal-to-noise ratio (SNR) conditions, none of them exhibited noticeable performance changes as the SNR decreased. Across all models, accuracy varied by only a few percentage points when moving from clean audio to severely degraded (1515dB) conditions.
Importantly, this observation should be interpreted within the specific context of auditory motion perception studied in AMPBench. The near-constant accuracy across SNR levels does not necessarily indicate general robustness to acoustic degradation; rather, it suggests that current LALMs fail to effectively leverage spatial cues whose reliability is expected to deteriorate under low-SNR conditions. As a result, reducing SNR does not further impair performance because the models already underutilize fine-grained spatial information embedded in the waveform.
This behavior contrasts sharply with human auditory perception, where motion discrimination performance degrades substantially as spatial cues such as ITD and ILD become unreliable in noisy environments.
Task-wise and Model-wise Differences.
Across tasks, we observe that Task 2, which involves variable-speed motion, is slightly easier for most models than Task 1 and Task 3, indicating that some forms of variation may be easier to capture than front–back ambiguities. Model-wise, Qwen2-Audio-7B tends to dominate in Task 1 and Task 2 at higher SNRs, while Voxtral-Mini-3B-2507 occasionally achieves the best scores at clean condition, reflecting different robustness trade-offs between architectures. Nevertheless, the overall performance remains far from satisfactory in all regimes.
Task-format-specific Evaluation.
@@
Table 4: MCQ vs T/F accuracy on AudioMotionBench (clean). Each value reports the mean accuracy with its standard deviation across repeated runs. Bold numbers indicate the best-performing model within each task and question format.
@@ 4.3 Prompt Ablation
To examine whether audio–language models genuinely verify auditory motion rather than following linguistic heuristics, we focus on the True/False (T/F) format of AudioMotionBench. Each audio sample is paired with two statements describing a motion trajectory. The True statement corresponds to the ground-truth direction, whereas the False statement is generated by mirroring the trajectory (e.g., left-to-right versus right-to-left) while matching lexical length, tense, and sentiment to eliminate textual shortcuts, as illustrated in Figure 3.
Figure 3: True/False prompt construction. Each audio sample is paired with a True statement describing the ground-truth motion and a False statement created by mirroring the trajectory while matching lexical length and phrasing.
@@
Table 5 summarizes the results under the clean condition of AudioMotionBench.Across all evaluated models, a consistent behavioral pattern appears within the scope of our motion–verification setting. The hit rate on True statements (TPR) is notably low, with the strongest model, Aero-1-Audio, accepting fewer than $21\%$ of True statements. In contrast, the rejection rate for False statements (TNR) is considerably higher, with Kimi-Audio-7B exceeding $84\%$. This imbalance suggests that under our evaluation protocol, models tend to be conservative when deciding whether a statement matches the underlying motion.
@@
Table 5: True/False verification on AudioMotionBench (clean). We report true positive rate (TPR), true negative rate (TNR), and YesBias (%), where lower YesBias indicates reduced tendency toward “Yes” responses. Bold values mark the best performance for each metric.
@@ 5 Conclusion
Summary of Findings.
In our work, we introduce AudioMotionBench, a physically grounded benchmark designed to systematically evaluate auditory motion perception in large audio–language models (LALMs). Through controlled experiments and cross-model comparisons, we reveal a consistent deficiency in spatial reasoning: while current models can often recognize what is heard, they still struggle to perceive where the sound source is and how it moves. This gap appears across architectures of different scales and training regimes, suggesting a structural limitation rather than model-specific behavior.
Key Insights.
@@
Broader Implications.
Beyond serving as a diagnostic benchmark, AudioMotionBench provides a principled foundation for studying spatially grounded auditory intelligence. The ability to track moving sound sources is essential for real-world, embodied agents that must locate, follow, or respond to dynamic auditory events in their environment. AudioMotionBench thus highlights the need for richer training data, architectural elements that explicitly model spatial cues, and learning objectives that reward accurate reasoning about physical motion. We hope this work encourages further exploration toward audio–language systems capable of perceiving the world not only semantically but also spatially, enabling a more embodied form of auditory understanding.
@@ 6 Limitations
While AudioMotionBench offers a controlled and fully reproducible testbed for studying auditory motion perception, it inevitably simplifies several aspects of real-world acoustic experience. The benchmark relies on analytically rendered spatial audio and a restricted set of motion patterns and sound types, which allows precise isolation of spatial cues but does not fully reflect the complexity of natural acoustic environments. In practice, factors such as reverberation, occlusion, interference, and diverse sound semantics may interact with motion perception in ways not captured by the current design.
As a result, performance on AudioMotionBench should be interpreted as a measure of a model’s ability to reason about idealized spatial motion cues, rather than a comprehensive assessment of real-world auditory intelligence. Extending the benchmark toward more realistic recordings, continuous motion dynamics, and richer sound categories remains an important direction for future work, building upon AudioMotionBench as an initial foundation.